# Postmerger: a new and dominant contribution to the gravitational-wave background from binary neutron stars

### v1

@@ Appendix B Modelling the BNS population
Here the SFR is taken from Ref. [72] and $P(t_{\mathrm{d}})$ is the distribution of time intervals $t_{\mathrm{d}}$ between the formation of stellar progenitors and the eventual BNS merger. We here assume a power-law distribution $P(t_{\mathrm{d}})\propto t_{\mathrm{d}}^{-1}$ [73], and the delay times to range from $t_{\mathrm{d},\mathrm{min}}=10\,{\rm Myr}$ to $t_{\mathrm{d},\mathrm{max}}$ set by the Hubble time. The parameter $R_{0}$ in Eq. (10) is a normalization constant that calibrates the merger rate at $z=0$ the local BNS merger rate inferred from the GWTC-3 LIGO-Virgo catalogue: $R_{0}=44^{+96}_{-34}~{}~{}{\rm yr}^{-1}\,{\rm Gpc}^{-3}$ [12].
Appendix C Network sensitivity for SGWB
The SGWB is in noise-dominated mode for ground-based detectors, as it is always much weaker than the detector noise. Moreover, its stochastic nature makes the use of match-filtering techniques, which have been developed and are effective for individual sources, inadequate. Instead, detection methods rely on cross-correlations of the outputs of pairs of detectors and searches for excess signal, under the assumption that the noise is only weakly correlated [101, 102]. Knowledge of the expected signal spectrum can help in constructing adapted searches, and if the spectrum is broad, integration over frequencies can increase the SNR. In order to better visualise the detectability of the SGWB, we use power-integrated (PI) sensitivity curves, following Ref. [103] and briefly summarised below. We also outline our calculation of PI sensitivity curves for a network of ET and CE detectors with different configurations.
We start by assuming the background has a power-law spectrum with slope $\sigma$ and amplitude $\Omega_{0}$
$\Omega_{\rm{GW}}(f)=\Omega_{0}\left(\frac{f}{f_{\rm{ref}}}\right)^{\alpha}\,.$ (11)
It can then be shown that $\alpha=2/3$ for the inspiral part of the spectrum. We further define the PSD $S_{h}(f)$ of the GW signal as [66]
$S_{h}(f):=\frac{3H_{0}^{2}}{2\pi^{2}}\frac{\Omega_{\rm{GW}}(f)}{f^{3}}\,,$ (12)
where $H_{0}$ is the Hubble parameter, and we are using a single-sided spectrum.
The broadband SNR for a network cross-correlation search for an unplaced and isotropic stochastic background $\Omega_{\mathrm{GW}}(f)$, integrated over frequency in the range $[f_{\mathrm{min}},f_{\mathrm{max}}]$, is given by [103]
$\rho:=\sqrt{2\tau}\left[\int^{f_{\mathrm{max}}}_{f_{\mathrm{min}}}\sum_{i=1}^{N}\sum_{j>i}^{N}\frac{\gamma^{2}_{ij}(f)S_{h}^{2}(f)\sin^{2}\beta_{i}\sin^{2}\beta_{j}}{P_{n,i}(f)P_{n,j}(f)}\mathrm{d}f\right]\,,$ (13)
where $\tau$ is the observation time, $\beta_{i}$ and $\beta_{j}$ are the opening angles of interferometers $i$ and $j$, respectively, $P_{n,i}$ and $P_{n,j}$ are their noise PSDs, and $\gamma(f)$ is the overlap reduction function that quantifies the correlation between the signals observed by the different interferometers [104].
It is important to note that for detectors that are located far apart or have large relative orientation angles, $\gamma(f)$ becomes an oscillatory function that approaches zero at high frequencies. This behaviour reduces the sensitivity of the detector pair in the frequency range of interest for this study (i.e., above $\sim 1\,{\rm kHz}$). Conversely, for co-located detectors, $\gamma(f)$ remains constant across all frequencies and varies only with the orientation angle, reaching a maximum value of $\gamma=1$ for co-aligned L-shaped detectors.
The PI sensitivity curve, constructed for a given SNR $\rho_{0}$ and observation time $\tau$, is defined such that a power-law SGWB tangent to this curve is detected with this SNR for an observation time $\tau$. More specifically, for a given slope $\alpha$ of the SGWB and for a given observation time $\tau$, we can find an amplitude $\Omega_{0,\alpha}$ such that the SNR in Eq. (13) is given by some fixed value $\rho=\rho_{0}$
$\Omega_{0,\alpha}:=\frac{\rho_{0}}{\sqrt{2\tau}}\,\left[\int^{f_{\mathrm{max}}}_{f_{\mathrm{min}}}\frac{(f/f_{\mathrm{ref}})^{2\alpha}}{\Omega_{\mathrm{eff}}^{2}}\mathrm{d}f\right]^{-1/2}\,.$ (14)
lat () lon () az () $\beta$ () $L({\rm km})$
${\rm ET}$ $40.5$ 9.49.4~{}~{}~{}9.49.4 $70.6$ $60$ $10$
${\rm CE}_{1}$ $46.5$ $-119.4$ 126126126~{}126 $90$ $40$
${\rm CE}_{2}$ $46.5$ $-119.4$ 126126126~{}126 $90$ $20$
Table 1: Latitude, longitude, azimuth, opening angle and arms length of the interferometers considered in this study. The corresponding noise PSDs are taken from Refs. [31, 105].
The effective background $\Omega_{\mathrm{eff}}$ is defined using Eq. (12) and the effective-strain noise PSD $S_{\mathrm{eff}}$ as [81]
$S_{\mathrm{eff}}(f):=\left[\sum_{i=1}^{N}\sum_{j>i}^{N}\frac{\gamma^{2}_{ij}(f)\sin^{2}\beta_{i}\sin^{2}\beta_{j}}{P_{n,i}(f)P_{n,j}(f)}\right]^{-1/2}\,,$ (15)
where, for a single interferometer, Eq. (15) reduces to $S_{\mathrm{eff}}=\sqrt{P_{\mathrm{n}}}=S_{\mathrm{n}}$, as expected. The PI sensitivity curve is then given by the envelope of all the power-law functions $\Omega_{\rm{GW}}$ with $\Omega_{0,\alpha}$ and $\alpha$ related through Eq. (14). We refer the reader to Ref. [103] for further details on this method. In the following, all the sensitivity curves we report are PI sensitivity curves.
Two different scenarios have been considered in this work: a single triangular $10\,{\rm km}$ ET detector and a network composed of 2 ET detectors ${\rm ET}_{1,2}$ combined with two CE detectors ${\rm CE}_{1,2}$. The latter case includes two triangular $10\,{\rm km}$ ET and two CEs: a $20\,{\rm km}$ L-shaped ${\rm CE}_{1}$ and a $40\,{\rm km}$ L-shaped ${\rm CE}_{2}$. All the parameters for each scenario are detailed in Tab. 1, while the PSDs for ET and CE are taken from Refs. [31] and [105], respectively. We note that we neglect the impact of correlated noise, which may however be significant for co-located ET detectors [31, 106]. Finally, we optimistically assume the detectors to operate at a duty cycle of $100\%$, and explore two possibilities for the observations time: $1$ and $10$ years.

### v2

@@ Abstract
We assume that the BNS merger rate follows the star formation rate (SFR) convolved with a distribution of delay times. Specifically, we adopt the baseline_delays model from Ref. Lehoucq et al. (2023), where the merger rate is defined as: $R_{\mathrm{merg}}(t)=R_{0}\int^{t_{\mathrm{d},\mathrm{max}}}_{t_{\mathrm{d},\mathrm{min}}}\mathrm{SFR}(t-t_{\mathrm{d}})\,P(t_{\mathrm{d}})\,\mathrm{d}t_{\mathrm{d}}$. Here the SFR is taken from Ref. Vangioni et al. (2015) and $P(t_{\mathrm{d}})$ is the distribution of time intervals $t_{\mathrm{d}}$ between the formation of stellar progenitors and the eventual BNS merger. We here assume a power-law distribution $P(t_{\mathrm{d}})\propto t_{\mathrm{d}}^{-1}$ Chruslinska et al. (2018), and the delay times to range from $t_{\mathrm{d},\mathrm{min}}=10\,{\rm Myr}$ to $t_{\mathrm{d},\mathrm{max}}$ set by the Hubble time. The parameter $R_{0}$ is a normalization constant that calibrates the merger rate at $z=0$ to the local BNS merger rate inferred from the GWTC-3 LIGO-Virgo catalogue: $R_{0}=44^{+96}_{-34}~{}~{}{\rm yr}^{-1}\,{\rm Gpc}^{-3}$ Abbott, R. et al. (2023).